please answer all the question

1.1 Data that is recorded on a 1 – 5 rating scale (e.g. 1 = poor, 2 = fair, 3 = good, 4 = very good, 5 = excellent) to a question such as “Rate the level of service you received from your most recent experience with your bank” is:

A. nominal data

B. continuous data

C. interval data

D. ratio data

E. ordinal data

1.2 A positively skewed histogram means that:

A. there are a few extremely small data value in the data set

B. there are more positive values in the data set

C. there are a few extremely large data values in the data set

D. the mean is a more reliable than the median as the measure of central location

E. none of the above

1.3 Which one of the following descriptive statistical measures cannot be read off from an ogive (the graph of a relative cumulative frequency distribution)?

A. the mean

B. the lower quartile

C. the median

D. the upper quartile

E. none of the above

1.4 If n = 10 and p = 0.70, then the standard deviation of the binomial distribution is:

A. 0.07

B. 1.45

C. 7.00

D. 14.29

E. none of the above

please answer all the question
Answer #1

(1.1)

Correct option:

E.   ordinal data

Explanation: Data that is recorded on a 1 – 5 rating scale (e.g. 1 = poor, 2 = fair, 3 = good, 4 = very good, 5 = excellent) to a question such as “Rate the level of service you received from your most recent experience with your bank” is ordinal data because it can be ranked.

(1.2)

Correct option:

C.   there are a few extremely large data values in the data set

Explanation:

In a positively skewed histogram, most values are clustered around the left tail of the distribution. The right tail of the distribution is longer.

(1.3)

Correct option:

A   the mean

Explanation:

Mean cannot be read off from an ogive (the graph of a relative cumulative frequency distribution)

(1.4)

Correct option:

B.    1.45

Explanation:

\sigma =\sqrt{npq}=\sqrt{10\times 0.70\times 0.30}=1.45